OpenStudy (anonymous):

Find sin θ if cot θ = - 2 and cos θ < 0

OpenStudy (anonymous):

cosec^2(theta)=1+4=5 or sin(theta)=1/(sqrt5) (as per the question theta lies in the second quadrant hence sin(theta) is +ve)

OpenStudy (anonymous):

wouldnt sin be -5?

OpenStudy (anonymous):

sin(theta) can never be less than -1

OpenStudy (anonymous):

ok so it would be sqrt5/2 right?

OpenStudy (anonymous):

can u give reasons for ur answers

OpenStudy (anonymous):

for simplicity sin(theta)=1/(cosec(theta))

OpenStudy (anonymous):

cosec(theta)=sqrt(1+cot^2(theta))

OpenStudy (anonymous):

ahh Thank you (: i understand now